An organic compound weighing x gm containing 46.67% of nitrogen, is heated with NaOH. It gave NH3 which is neutralised by 200ml of 1N HCl. Find x.


Grade:12

1 Answers

Arun
3 years ago
let the organic compund be represented by X
x +NaOH ------NH3 +HCL----NH4Cl
mole of NH3 formed is 200*1/1000=.2mole
weight of nitrogen in NH3=.2*14=2.8gm
therefore weight in NH3 and weight of nitrogen in organic compound will be same
thus the percentage of nitrogen in organic compound =(2.8/6)*100=46.67